OpenStudy (anonymous):

I AM REVISING -THE DEEPER I GO, THE HARDER! INTEGRATE x(2^x)

OpenStudy (anonymous):

\[\int\limits_{}^{}x2^{x}\]

OpenStudy (owlcoffee):

When we have the product of functions in a integral, it's a little hard to solve, because I can't make any substitution or direct solving. But remembering one thing the derivative of a product, it had this form: \[(g(x)f(x))'=g'(x)f(x)+g(x)f'(x)\] now, something what happens if I integrate these?: \[\int\limits_{}^{} (g(x)f(x))' dx=\int\limits g'(x)f(x) dx + \int\limits g(x)f'(x) dx \] I'll re-write these: \[\int\limits g(x)f'(x) dx = \int\limits (g(x)f(x))' dx - \int\limits g'(x)f(x) dx\] Thre is a very interesting theorem that states this: \[\int\limits u' dx = u\] where u is a function and u' it's derivative form. And also: \[g'(x)=\frac{ dg }{ dx } => g'(x)dx=dg\] \[f'(x)=\frac{ df }{ dx } => f'(x)dx=df\] Then, replacing: \[\int\limits g(x)df=g(x)f(x)- \int\limits f(x)dg\] you might know this formula this way: \[\int\limits u dv = uv - \int\limits v du\] And that is what we call "integral by parts" formula.

OpenStudy (anonymous):

ok, let me try to follow here

OpenStudy (owlcoffee):

First try choosing wich one will be "u" and wich one will be "v".

OpenStudy (owlcoffee):

there is a good way of knowing wich one to choose. it's called ILATE, or in other words: "Inverse, Logarithmic, Algebraic, Trigonometric, Exponential". "x" is algebraic and "2exp x " is exponential. So it's conveninient to choose "x" as "u" and "2exp x " as "v"

OpenStudy (anonymous):

oh, thank for ILATE

OpenStudy (owlcoffee):

Okay, you can show me how you're doing when you try it out :)

OpenStudy (anonymous):

YES

OpenStudy (nincompoop):

u =x, u' = dx v' = 2^x dx, v = (2^x)/log 2

OpenStudy (anonymous):

BUT THEN it takes me back to \[\int\limits_{}^{} x2^{x}=\]\[x \frac{ 2^{x} }{ \log2 }- \int\limits_{}^{}x\frac{ 2^{x} }{ \log2 }\]

OpenStudy (anonymous):

@owlcoffee

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

i am stuck, seriously...

hartnn (hartnn):

how did u get 'x' inside 2nd integral? isn't du = dx ? and not x \(\large x \frac{ 2^{x} }{ \log2 }- \int\limits_{}^{}\frac{ 2^{x} }{ \log2 }\color{green}{dx}\)

OpenStudy (owlcoffee):

log2 is a constant, you can pull it out of the integral. and integrate 2 exp x. I'm getting very rusty in mathematics.

OpenStudy (anonymous):

oh, ok, let me correct that

OpenStudy (anonymous):

\[\int\limits_{}^{}x2^{x}dx= \frac{x2^{x}}{\log2} -\int\limits_{}^{}\frac{ 2^{x} }{ \log2 }dx\] = \[\frac{ x2^{x} }{ \log2 }-\frac{ 2^{x} }{( \log2)^{2} }\]

OpenStudy (anonymous):

i think this z ryt

OpenStudy (tkhunny):

It's By Parts, not magic. \(\int x\cdot 2^{x}\;dx = \int x\;d\left(\dfrac{2^{x}}{\log(2)}\right) = \dfrac{x\cdot 2^{x}}{\log(2)} - \int\dfrac{2^{x}}{\log(2)}\;dx\)

OpenStudy (amistre64):

by parts can be organized in a table: